OpenStudy (anonymous):

Solve the equation 25 = sqrt5^(2x + 6). Using complete sentences, explain the strategy used in solving this equation.

OpenStudy (anonymous):

Square both sides

OpenStudy (anonymous):

You need to use equation editor, otherwise I can't understand the equation!

OpenStudy (dcolley):

It looks like it should be something like this...I guess \[25=\sqrt{5^(2x+6}\] And if Im correct... then you can solve by using the property of logarithms

OpenStudy (anonymous):

\[25=\sqrt{5^{2x+6}}=(\sqrt{5})^{2x+6} \implies 2x+6=4 \implies x= -1\]